The graph of y=6 - [SQRT(4 - 3x)] contains which point?
A (1,4)
B(0,1)
C(0,6)
D(1,5)
E (2,4)
F none of the above



9y2 + 16=0
A no solution
B4/3
C-4/3
D{-4/3, 4/3}
E 16/9
F -1

Answer by Cromlix(4381)

y = 6 - [SQRT(4 - 3x)]
D: 1,5
y = 6 -[SQRT(4 - 3(1)]
y = 6 -[SQRT(1)]
y = 5.
...........
9y^2 + 16 = 0
9y^2 = -16
y^2 = -16/9
A
No solution as cannot take
square root of negative number..
Hope this helps.
:-)